OpenStudy (anonymous):

Help me please me Find the value of x and y a.)x=109,y=108 b.)x=108,y=109 c.)x=108,y=108 d.)x=109,y=109

OpenStudy (yamyam70):

the question is insufficient

OpenStudy (anonymous):

@yamyam70 man fuch yo spongebob Azz

OpenStudy (anonymous):

i have to post the picture

OpenStudy (anonymous):

give me a second okay

OpenStudy (yamyam70):

alright

OpenStudy (chosenmatt):

need help still?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes indeed of help

OpenStudy (shaadykhamlichi):

y=109 and x=108

OpenStudy (shaadykhamlichi):

and i'm sure about that

OpenStudy (yamyam70):

how did you get x ?

OpenStudy (anonymous):

alrighty thank you so much :)

OpenStudy (shaadykhamlichi):

all shapes that has 4 sides the total of the angle is 360 so: 360-(100+81+71)=108

OpenStudy (shaadykhamlichi):

@yamyam70

OpenStudy (yamyam70):

oh thanks :) and solving for y is .. -180 right?

OpenStudy (shaadykhamlichi):

yes y=180-71=109

OpenStudy (yamyam70):

Thanks mate : )